OpenStudy (theislander):

What is the standard form of the equation of a circle that has its center at (-2, -3) and passes through the point (-2, 0)?

OpenStudy (campbell_st):

well the distance from the centre to the point is the radius the standard form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre and r is the radius... you know h = -2 and k = -3 and you can find r... so substitute and perhaps expand... hope this helps

OpenStudy (theislander):

would it be (x − 2)2 + (y − 3)2 = 16 or (x − 2)2 + (y − 3)2 = 4 @campbell_st

OpenStudy (campbell_st):

the distance from the centre (-2, -3) to (-2, 0) is 3 units up so the equation is (x + 2)^2 + (y + 3)^2 = 3^2